# ASVspoof 5: Design, Collection and Validation of Resources for Spoofing, Deepfake, and Adversarial Attack Detection Using Crowdsourced Speech

### v3

@@ 5 Visualisation
Illustrated in Figure 6 is a representation of the ASVspoof 5 database in the form of a 2-dimensional t-SNE plot [107]. Each point corresponds to an embedding extracted using an ASV system (the baseline system described in § 7). Gray points correspond to bona fide utterances while coloured points correspond to one of the 32 attacks. To reduce computation time and visual clutter, points are shown for a random selection of 10% of the utterances in the training, development and evaluation sets, with and without encoding/compression for the evaluation set only. The recipe used to produce Figure 6 is the same as that used to produce similar plots in [3] and [9] and involves attack-level, within-class covariance normalisation, length normalisation and speaker-level data whitening. Also plotted as a gray solid line in Figure 6 is a confidence contour which encloses 98.9% of the probability mass of the bona fide data distribution. It is estimated using a kernel density estimate (KDE) method. [^18] We use the Seaborn KDE API [109]. The confidence level of 98.9% was used in deriving a similar visualisation for the ASVspoof 2019 database [9, Figure 3], though in the form of a confidence ellipse.
@@
The dendrogram (and the similarity heatmap) show similar clusters of attacks as the t-SNE plot, e.g., {A01, A02, A03} and {A04, A05, A06}. The high similarities within these two groups of TTS systems may be because the three systems within each group use the same acoustic decoder, text encoder and vocder, even if they use different speaker embedding extractors. However, the two sub-groups have low inter-group similarity – they are located in different branches of the dendrogram tree. We observe a similar clustering of {A09, A10, A21, A22}, all implemented using the ToucanTTS toolkit.
@@ 6 TTS and VC system optimisation using surrogate models
Figure 8 illustrates the increase in ASV and CM equal error rates (EERs) which three data contributors (of five in total) were able to achieve using ASV and CM surrogates. The four plots show EERs for each of the four surrogate models (1 ASV and 3 CM systems) and for a selection of three attacks {A10, A11, A24} the performance of which was evaluated using the surrogate models in multiple rounds (horizontal axis). The contributors decided the number of rounds and received results for their own attacks after each around. A10 contributors monitored the training process and selected the checkpoint that achieved the highest ASV EER. A11 contributors selected the checkpoint that achieved the highest overall surrogate CM EER. A24 contributors manually tweaked the training configuration of the speaker encoder (e.g. from independent training to joint training with the VC system) and selected the checkpoint that gave the highest surrogate ASV EER.
@@
Figure 8: Progress in ASV and CM system optimisation using surrogate models.
@@ 9 Acknowledgement
This work is partially supported by JST, PRESTO Grant Number JPMJPR23P9, Japan, and with funding received from the French Agence Nationale de la Recherche (ANR) via the BRUEL (ANR-22-CE39-0009) and COMPROMIS (ANR22-PECY-0011) projects. This work was also partially supported by the Academy of Finland (Decision No. 349605, project “SPEECHFAKES”). Part of the computation and data generation is done out using the TSUBAME4.0 supercomputer at Tokyo Institute of Science, and by supercomputing infrastructure provided by CSC — IT Center for Science (Finland).
@@ Appendix B Dendrogram of ASVspoof 5 data
where ||$\cdot$|| ⋅ | denotes the number of observations and $c(\boldsymbol{a},\boldsymbol{b})$ is cosine distance of vectors $\boldsymbol{a}$ and $\boldsymbol{b}$. The similarity matrix $\boldsymbol{S}$ is used for hierarchical agglomerative clustering, using the Ward’s minimum variance criterion as the clustering objective. The figure is rendered using the Seanborn clustermap API [109].

### v4

@@ 5 Visualisation
Illustrated in Figure 6 is a representation of the ASVspoof 5 database in the form of a 2-dimensional t-SNE plot [107]. Each point corresponds to an embedding extracted using an ASV system (the baseline system described in § 7). Grey points correspond to bona fide utterances while coloured points correspond to one of the 32 attacks. To reduce computation time and visual clutter, points are shown for a random selection of 10% of the utterances in the training, development and evaluation sets, with and without encoding/compression for the evaluation set only. The recipe used to produce Figure 6 is the same as that used to produce similar plots in [3] and [9] and involves attack-level, within-class covariance normalisation, length normalisation and speaker-level data whitening. Also plotted as a grey solid line in Figure 6 is a confidence contour which encloses 98.9% of the probability mass of the bona fide data distribution. It is estimated using a kernel density estimate (KDE) method. [^18] We use the Seaborn KDE API [109]. The confidence level of 98.9% was used in deriving a similar visualisation for the ASVspoof 2019 database [9, Figure 3], though in the form of a confidence ellipse.
@@
The dendrogram (and the similarity heatmap) show similar clusters of attacks as the t-SNE plot, e.g., {A01, A02, A03} and {A04, A05, A06}. The high similarities within these two groups of TTS systems may be because the three systems within each group use the same acoustic decoder, text encoder and vocoder, even if they use different speaker embedding extractors. However, the two sub-groups have low inter-group similarity – they are located in different branches of the dendrogram tree. We observe a similar clustering of {A09, A10, A21, A22}, all implemented using the ToucanTTS toolkit.
@@ 6 TTS and VC system optimisation using surrogate models
Figure 8 illustrates the increase in ASV and CM equal error rates (EERs) which three data contributors (of five in total) were able to achieve using ASV and CM surrogates. The four plots show EERs for each of the four surrogate models (1 ASV and 3 CM systems) and for a selection of three attacks {A10, A11, A24} the performance of which was evaluated using the surrogate models in multiple rounds (horizontal axis). The contributors decided the number of rounds and received results for their own attacks after each round. A10 contributors monitored the training process and selected the checkpoint that achieved the highest ASV EER. A11 contributors selected the checkpoint that achieved the highest overall surrogate CM EER. A24 contributors manually tweaked the training configuration of the speaker encoder (e.g. from independent training to joint training with the VC system) and selected the checkpoint that gave the highest surrogate ASV EER.
@@
Figure 8: Progress in TTS and VC system optimisation, measured using the EERs of the surrogate models.
@@ 9 Acknowledgement
This work is partially supported by JST, PRESTO Grant Number JPMJPR23P9, JST AIP Acceleration Project Grant Number JPMJCR24U3, Japan, and with funding received from the French Agence Nationale de la Recherche (ANR) via the BRUEL (ANR-22-CE39-0009) and COMPROMIS (ANR22-PECY-0011) projects. This work was also partially supported by the Academy of Finland (Decision No. 349605, project “SPEECHFAKES”). Part of the computation and data generation was performed using the TSUBAME4.0 supercomputer at Tokyo Institute of Science, and by supercomputing infrastructure provided by CSC — IT Center for Science (Finland).
@@ Appendix B Dendrogram of ASVspoof 5 data
where ||$\cdot$|| ⋅ | denotes the number of observations and $c(\boldsymbol{a},\boldsymbol{b})$ is cosine distance of vectors $\boldsymbol{a}$ and $\boldsymbol{b}$. The similarity matrix $\boldsymbol{S}$ is used for hierarchical agglomerative clustering, using the Ward’s minimum variance criterion as the clustering objective. The figure is rendered using the Seaborn clustermap API [109].